# Quantum Algorithms in a Superposition of Spacetimes

### v2

@@ 3.2 Statistical and Algorithmic Tools
There exists a quantum algorithm $Q_{ST}\left(\cdot\right)$ that gets as input a $2\cdot n$-qubit register $R$ and outputs a bit $b\in\{0,1\}$, and executes in complexity $O(n)$. If $R$ is in the separable state of $n$ qubits each $|{\psi}\rangle\otimes|{\phi}\rangle$, then the algorithm outputs $1$ with probability $\frac{1}{2}+\frac{|\langle{\psi}||{\phi}\rangle|^{2}}{2}$, and with the remaining probability outputs 00.
@@
As part of the reduction $\mathsf{GapCVP}\leq_{p}\mathsf{SISD}_{\mathsf{poly},1}$, which is resented in Section 7, we will approximate the discrete truncated Gaussian distribution, which is defined as follows.
@@
$\forall x\in\mathbb{Z}:\mathcal{N}^{\mathbb{Z}}\left(\mu,\sigma\right)\left(x\right):=\int_{x-\frac{1}{2}}^{x+\frac{1}{2}}\mathcal{N}\left(\mu,\sigma\right)\left(x\right)\enspace,$
@@
We use the following fact on the total variation distance between the two versions (discrete and rounded) of the truncated Gaussian distributions.
Fact 3.3 (Distance between rounded and discrete Gaussians).
Let $B\in\mathbb{N}$, then,
$\left\lVert\left(\mathcal{N}^{B}\left(0,B\right)\right)-\left(D_{B}\right)\right\rVert_{TV}\leq O\left(\frac{1}{B}\right)\enspace.$
@@
The central limit theorem (CLT), one of the fundamental theorems in statistics, tells us that when we take $N\in\mathbb{N}$ independent and continuous random variables $X_{1},\cdots,X_{N}$, the uniform average of the samples approaches the Gaussian distribution, as $N$ goes to infinity. The CLT does not tell us the rate at which the samples approach the Gaussian distribution, as this depends, among other things, on the probability density functions of the summed variables $\{X_{i}\}_{i\in[N]}$. The Berry-Esseen Theorem is a quantified version of the central limit theorem – it bounds the Wasserstein-1 distance between the sum $S:=\sum_{i\in[N]}X_{i}$ and the Gaussian distribution, as a function of $N$ and the density functions of the variables $\{X_{i}\}_{i\in[N]}$. The original formulation of the Berry-Essen Theorem is insufficient for our needs, because (1) In this paper we are working with discrete random variables and not continuous, and (2) We want a bound on the total variation distance which implies physical statistical indistinguishability, and for this, the Wasserstein-1 distance is not enough. We will use the following strengthening Theorem from [CGS10]:
Theorem 3.3 (Discrete Berry-Esseen bound for total variation – Theorem 7.4 from [CGS10]).
Let $N\in\mathbb{N}$ and let $\{X_{i}\}_{i\in[N]}$ independent random variables with finite supports contained in $\mathbb{Z}$, and denote $S:=\sum_{i\in[N]}X_{i}$. Denote,
  • For every $i\in[N]$, the expectation of $X_{i}$, $\mu_{i}:=\mathbb{E}\left(X_{i}\right)$.
  • For every $i\in[N]$, the variance of $X_{i}$, $\sigma_{i}^{2}:=\text{Var}\left(X_{i}\right):=\mathbb{E}\left(X_{i}^{2}\right)-\mathbb{E}\left(X_{i}\right)^{2}$.
  • The variance of the sum $S$, $\sigma^{2}:=\text{Var}\left(S\right)$, which equals $\sum_{i\in[N]}\sigma_{i}^{2}$, due to the random variables being independent.
  • For every $i\in[N]$, the third moment of $X_{i}$, $\gamma_{i}:=\frac{\mathbb{E}_{X_{i}}\left(|X_{i}-\mu_{i}|^{3}\right)}{\sigma^{3}}$.
  • The expectation of $S$, $\mu:=\sum_{i\in[N]}\mu_{i}$.
  • The sum of third moments $\gamma:=\sum_{i\in[N]}\gamma_{i}$.
  • The sum of the maximal values that the variables can take, $B:=\sum_{i\in[N]}\left(\max_{x\in X_{i}}|x|\right)$.
  • For every $i\in[N]$, the random variable $S^{(i)}:=S-X_{i}$, that is, the original full sum $S$, excluding the $i$-th variable.
@@
$\left\lVert S-\mathcal{N}^{B}\left(\mu,\sigma\right)\right\rVert_{TV}$
$\leq\frac{3}{2}\cdot\sigma\cdot\sum_{i\in[N]}\left(\left(\gamma_{i}+\frac{2\cdot\sigma_{i}^{2}}{3\cdot\sigma^{3}}\right)\cdot\left\lVert\left(S^{(i)}\right)-\left(S^{(i)}+1\right)\right\rVert_{TV}\right)$
@@ 4.1 Quantum Computation with Computable Order Interference
Remark 4.2 (Stronger versions of the OI oracle).
@@ 6 $\mathsf{GI}\leq_{p}\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
7 $\mathsf{GapCVP}_{O\left(n^{2}\right)}\leq_{p}\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
@@
$\mathsf{GapCVP}_{c\cdot n^{2}}\leq_{p}\mathsf{SISD}_{\left(a(n),b(n),r(n)\right)}\enspace,$
for $a(n):=\frac{1}{2}+\frac{1}{2^{n}}$, $b(n):=1$, $r(n):=1$.
@@
Let $\left(\mathbf{B},\mathbf{t},d\right)$ an input for the $\mathsf{GapCVP}_{g(n)}$ problem, for $g(n)=c_{g}\cdot n^{2}$ where $c_{g}:=256\cdot\frac{c_{0}}{c_{1}}\cdot\pi$, where $c_{0},c_{1}\in\mathbb{R}_{>0}$ are two positive absolute constants we will specify later, $n\in\mathbb{N}$ is the dimension of the lattice basis $\mathbf{B}$ and the vector $\mathbf{t}$, and $b\in\mathbb{N}$ is the size of the binary representation of the largest (in absolute value) number in $\mathbf{B}$ and the vector $\mathbf{t}$. So, the description size of the input is
@@
  • $\beta:=\lfloor\log_{2}\left(B\right)\rfloor$, for $B:=\frac{g(n)\cdot d}{2\cdot\kappa\cdot\sqrt{n}}$.
@@
by using the first bit as a decider for the sign of the number, and the rest of the $s$ bits as the binary representation of the absolute value of the number (the representation is not necessarily bijective, of course). In the following proof we thus use binary strings of size $s+1$ to represent numbers in the set $\mathbb{Z}_{\left(-\left(2^{s}-1\right),2^{s}-1\right)}$.
@@
We define an invertible circuit sequence $C:=\left(C_{i,\rightarrow},C_{i,\leftarrow}\right)_{i\in[\ell]}$ below, where all $\ell$ circuit pairs have input/output size $n\cdot\left(s+1\right)$ (which we think of as an integer vector $\mathbf{v}\in\mathbb{Z}^{n}_{\left(-\left(2^{s}-1\right),2^{s}-1\right)}$) and randomness bounded by $1$ (that is, each circuit either uses a single bit of randomness or it is deterministic, and does not use randomness at all). In the below circuits, all arithmetic (which comes down to additions and multiplications) is modulo in absolute value of $12^{s}-$ 1.
@@
  • The output distribution of $C^{0}$ is the variable $\left(\mathbf{v}+\mathbf{e}\right)\in\mathbb{Z}^{n}_{\left(-\left(2^{m+b}\cdot n+\kappa\cdot 2^{\beta}\right),2^{m+b}\cdot n+\kappa\cdot 2^{\beta}\right)}$, where $\mathbf{v}\in\mathbb{Z}^{n}_{\left(-\left(2^{m+b}\cdot n\right),2^{m+b}\cdot n\right)}$ is a random lattice vector with a uniformly random coordinates vector $\mathbf{a}\in\mathbb{Z}^{n}_{\left(-2^{m},2^{m}-1\right)}$, and $\mathbf{e}\in\mathbb{Z}^{n}_{\left(-\kappa\cdot 2^{\beta},\kappa\cdot 2^{\beta}\right)}$ such that for each $j\in[n]$, the coordinate $\mathbf{e}_{j}$ is a random variable defined as follows. Since we subtracted at the end $\frac{\kappa}{2}\cdot\left(2^{\beta+1}-1\right)$, it is equivalent to subtracting $2^{\beta}$, $\frac{\kappa}{2}$ times followed by subtracting $\left(2^{\beta}-1\right)$, $\frac{\kappa}{2}$ times. We can thus consider the variable $\mathbf{e}_{j}$ as the sum of $\kappa/2$ i.i.d. uniformly random samples from $\mathbb{Z}_{\left(-2^{\beta},2^{\beta}-1\right)}$ and $\kappa/2$ i.i.d. uniformly random samples from $\mathbb{Z}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}$.

### v3

@@ 3.2 Statistical and Algorithmic Tools
There exists a quantum algorithm $Q_{ST}\left(\cdot\right)$ that gets as input a $2n$-qubit register $R$ and outputs a bit $b\in\{0,1\}$, and executes in complexity $O(n)$. If there exist two $n$-qubit states $|{\psi}\rangle$, $|{\phi}\rangle$ such that $R$ is in the separable state $|{\psi}\rangle\otimes|{\phi}\rangle$, then the algorithm outputs $1$ with probability $\frac{1}{2}+\frac{|\langle{\psi}||{\phi}\rangle|^{2}}{2}$, and with the remaining probability outputs 00.
@@
As part of the reduction $\mathsf{GapCVP}\leq_{p}\mathsf{SISD}_{\mathsf{poly},1}$, which is presented in Section 7, we will approximate the discrete truncated Gaussian distribution, which is defined as follows.
@@
$\forall x\in\mathbb{Z}:\mathcal{N}^{\mathbb{Z}}\left(\mu,\sigma\right)\left(x\right):=\int_{(x-\frac{1}{2},x+\frac{1}{2}]}\mathcal{N}\left(\mu,\sigma\right)\left(x\right)\enspace,$
@@
We use the following known statistical relations between the rounded Gaussian distribution and variations of the Gaussian distribution.
Fact 3.3 (Distance properties of the rounded Gaussian).
There exists a positive absolute constant $c_{G}\in\mathbb{R}_{>0}$ such that, for every $B\in\mathbb{N}$, all of the above total variation distances are bounded by $\frac{c_{G}}{B}$:
$\left\lVert\mathcal{D}\left(\mathcal{N}^{\mathbb{Z}}\left(0,B-1\right)\right)-\mathcal{D}\left(\mathcal{N}^{\mathbb{Z}}\left(0,B\right)\right)\right\rVert_{TV}\enspace,$
$\left\lVert\mathcal{D}\left(\mathcal{N}^{\mathbb{Z}}\left(0,B\right)\right)-\mathcal{D}\left(\mathcal{N}^{B}\left(0,B\right)\right)\right\rVert_{TV}\enspace,$
$\left\lVert\mathcal{D}\left(\mathcal{N}^{B}\left(0,B\right)\right)-\mathcal{D}\left(D_{B}\right)\right\rVert_{TV}\enspace.$
Fact 3.4 (Multivariate discrete Gaussian vector norm concentration bounds).
There exists an absolute positive constant $c_{D}\in\mathbb{R}_{>0}$ such that for all $B,n\in\mathbb{N}$,
$\Pr_{u\leftarrow D_{B}^{n}}\Big{[}\frac{\sqrt{n}\cdot B}{\sqrt{2}}\leq\left\lVert u\right\rVert\leq\left(2-\frac{1}{\sqrt{2}}\right)\cdot\sqrt{n}\cdot B\Big{]}\geq 1-e^{-c_{D}\cdot n}\enspace.$
We also use the following standard property of total variation distance between functions of independent r.v.’s.
Fact 3.5 (Total variation distance between switched-input random variable).
Let $X,Y$ independent random variables and let $f,g$ (not necessarily deterministic) functions on the union of the supports of $X$ and $Y$. Then, there exists a number $d_{\left(X,Y\right)}\in[-\left\lVert\mathcal{D}\left(X\right)-\mathcal{D}\left(Y\right)\right\rVert_{TV},\left\lVert\mathcal{D}\left(X\right)-\mathcal{D}\left(Y\right)\right\rVert_{TV}]$, such that
$\left\lVert\mathcal{D}\left(f\left(X\right)\right)-\mathcal{D}\left(g\left(X\right)\right)\right\rVert_{TV}=\left\lVert\mathcal{D}\left(f\left(Y\right)\right)-\mathcal{D}\left(g\left(Y\right)\right)\right\rVert_{TV}+2\cdot d_{\left(X,Y\right)}\enspace.$
@@
The central limit theorem (CLT), one of the fundamental theorems in statistics, tells us that when we take $\kappa\in\mathbb{N}$ independent and continuous random variables $X_{1},\cdots,X_{\kappa}$ with mean 00, the random variable $\frac{\sum_{i\in[\kappa]}X_{i}}{\sqrt{\kappa}}$ approaches the Gaussian distribution, as $\kappa$ goes to infinity. Our main barrier with using the CLT is the combination of two facts: (1) The CLT needs division for convergence, i.e. it looks at the variable $\frac{\sum_{i\in[\kappa]}X_{i}}{\sqrt{\kappa}}$ rather than $\sum_{i\in[\kappa]}X_{i}$, and (2) As part of the Technical Overview (Section 2) and Section 7 we explain why we need to sample from the Gaussian distribution using a sequence of invertible functions. As we are dealing with integers only, this means that in order to use the CLT we need to execute integer (i.e., rounded) division, which is not invertible. Other, more minor problems are that the CLT guarantees convergence in Wasserstein-1 distance and we would like convergence in the stronger total variation distance, and finally, the CLT does not tell us the rate at which the samples approach the Gaussian distribution, as this depends, among other things, on the probability density functions of the summed variables $\{X_{i}\}_{i\in[\kappa]}$. We will use the following strengthening Theorem from [CGS10]:
Theorem 3.3 (Discrete un-normalized Berry-Esseen bound for total variation distance – Theorem 7.4 from [CGS10]).
Let $\kappa\in\mathbb{N}$ and let $\{X_{i}\}_{i\in[\kappa]}$ independent random variables with finite supports contained in $\mathbb{Z}$, and denote $S:=\sum_{i\in[\kappa]}X_{i}$. Denote,
  • For every $i\in[\kappa]$, the expectation of $X_{i}$, $\mu_{i}:=\mathbb{E}\left(X_{i}\right)$.
  • For every $i\in[\kappa]$, the variance of $X_{i}$, $\sigma_{i}^{2}:=\text{Var}\left(X_{i}\right):=\mathbb{E}\left(X_{i}^{2}\right)-\mathbb{E}\left(X_{i}\right)^{2}$.
  • The variance of the sum $S$, $\sigma^{2}:=\text{Var}\left(S\right)$, which equals $\sum_{i\in[\kappa]}\sigma_{i}^{2}$, due to the random variables being independent.
  • For every $i\in[\kappa]$, the third moment of $X_{i}$, $\gamma_{i}:=\frac{\mathbb{E}_{X_{i}}\left(|X_{i}-\mu_{i}|^{3}\right)}{\sigma^{3}}$.
  • The expectation of $S$, $\mu:=\sum_{i\in[\kappa]}\mu_{i}$.
  • The sum of third moments $\gamma:=\sum_{i\in[\kappa]}\gamma_{i}$.
  • For every $i\in[\kappa]$, the random variable $S^{(i)}:=S-X_{i}$, that is, the original full sum $S$, excluding the $i$-th variable.
@@
$\left\lVert\mathcal{D}\left(S\right)-\mathcal{D}\left(\mathcal{N}^{\mathbb{Z}}\left(\mu,\sigma\right)\right)\right\rVert_{TV}$
$\leq\frac{3}{2}\cdot\sigma\cdot\sum_{i\in[\kappa]}\left(\left(\gamma_{i}+\frac{2\cdot\sigma_{i}^{2}}{3\cdot\sigma^{3}}\right)\cdot\left\lVert\mathcal{D}\left(S^{(i)}\right)-\mathcal{D}\left(S^{(i)}+1\right)\right\rVert_{TV}\right)$
@@ 4.1 Quantum Computation with Computable Order Interference
Remark 4.2 (Different versions of the OI oracle).
@@ 6 $\mathsf{GI}\leq_{p}\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
7 $\mathsf{GapCVP}_{O\left(n\sqrt{n}\right)}\leq_{p}\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
@@
$\mathsf{GapCVP}_{c\cdot n\sqrt{n}}\leq_{p}\mathsf{SISD}_{\left(a(n),b(n),r(n)\right)}\enspace,$
for $a(n):=\frac{1}{4}+O\left(n^{-\frac{1}{4}}\right)$, $b(n):=\frac{3}{4}-e^{-\Omega\left(n\right)}$, $r(n):=1$.
@@
Let $\left(\mathbf{B},\mathbf{t},d\right)$ an input for the $\mathsf{GapCVP}_{g(n)}$ problem, for $g(n)=c_{g}\cdot n\sqrt{n}$ where $c_{g}:=64\cdot c_{0}$, where $c_{0}\in\mathbb{R}_{>0}$ is a positive absolute constant we will specify later, $n\in\mathbb{N}$ is the dimension of the lattice basis $\mathbf{B}$ and the vector $\mathbf{t}$, and $b\in\mathbb{N}$ is the size of the binary representation of the largest (in absolute value) number in $\mathbf{B}$ and the vector $\mathbf{t}$. So, the description size of the input is
@@
  • $\beta:=\lfloor\log_{2}\left(B\right)\rfloor$, for $B:=\frac{g(n)\cdot d}{c_{\mathcal{N}}\cdot\sqrt{n}\cdot\sqrt{\kappa}}$, where $c_{\mathcal{N}}:=8$. For the algorithm to work we will need $\beta\geq 0$, which in turn happens if $B\geq 1$, and indeed:
    $g(n):=64\cdot c_{0}\cdot n\sqrt{n}:=8\cdot\sqrt{\kappa}\sqrt{n}:=c_{\mathcal{N}}\cdot\sqrt{\kappa}\sqrt{n}\enspace,$
    which, since $d\in\mathbb{N}$, implies $B\geq 1$.
@@
by using the first bit as a decider for the sign of the number, and the rest of the $s$ bits as the binary representation of the absolute value of the number (the only number that has two such binary representations, and not one representation, is 00). In the following proof we thus use binary strings of size $s+1$ to represent numbers in the set $\mathbb{Z}_{\left(-\left(2^{s}-1\right),2^{s}-1\right)}$.
@@
We define an invertible circuit sequence $C:=\left(C_{i,\rightarrow},C_{i,\leftarrow}\right)_{i\in[\ell]}$ below, where all $\ell$ circuit pairs have input/output size $n\cdot\left(s+1\right)$ (which we think of as an integer vector $\mathbf{v}\in\mathbb{Z}^{n}_{\left(-\left(2^{s}-1\right),2^{s}-1\right)}$) and randomness bounded by $1$ (that is, each circuit either uses a single bit of randomness, or it is deterministic and does not use randomness at all). In the below circuits, all arithmetic (which comes down to additions only) is modulo in absolute value of $12^{s}-$ 1, that is, once it gets out of bounds of one side (right, $12^{s}-$ 1 or left, $-\left(2^{s}-1\right)$), it pops up, continuing in the same direction, at the opposite side (left or right, respectively).
@@
  • The output distribution of $C^{0}$ is the variable $\left(\mathbf{v}+\mathbf{e}\right)\in\mathbb{Z}^{n}_{\left(-\left(2^{m+b}\cdot n+\kappa\cdot 2^{\beta}\right),2^{m+b}\cdot n+\kappa\cdot 2^{\beta}\right)}$, where $\mathbf{v}\in\mathbb{Z}^{n}_{\left(-\left(2^{m+b}\cdot n\right),2^{m+b}\cdot n\right)}$ is a random lattice vector with a uniformly random coordinates vector $\mathbf{a}\in\mathbb{Z}^{n}_{\left(-2^{m},2^{m}-1\right)}$, and $\mathbf{e}\in\mathbb{Z}^{n}_{\left(-\kappa\cdot 2^{\beta},\kappa\cdot 2^{\beta}\right)}$ such that for each $j\in[n]$, the coordinate $\mathbf{e}_{j}$ is a random variable defined as follows: We sample $\kappa$ i.i.d. samples from the distribution $\mathcal{U}_{\left(0,2^{\beta+1}-1\right)}$ (the discrete uniform distribution on the set $\{0,1,\cdots,2^{\beta+1}-1\}$), add all $\kappa$ samples together, and then subtract $\frac{\kappa}{2}\cdot\left(2^{\beta+1}-1\right)$, which is exactly the expectation of the sum, and is also an integer since $\kappa$ is an even natural number.
@@
To conclude the part where we describe the reduction and its properties, we have the equality
$\left\lVert\mathcal{D}\left(C^{0}\right)-\mathcal{D}\left(C^{1}\right)\right\rVert_{TV}$
$:=\left\lVert\mathcal{D}\left(\mathbf{v}+\mathbf{e}\right)-\mathcal{D}\left(\mathbf{v}+\mathbf{e}+\mathbf{t}\right)\right\rVert_{TV}\enspace.$
Next, let $\tilde{B}:=\lceil\sqrt{\kappa}\cdot\tilde{\sigma}\rceil$, $\tilde{\sigma}:=\sqrt{\frac{2^{2\beta+2}+2^{\beta+2}}{12}}$, and let $D^{n}_{\tilde{B}}$ the $n$-dimensional discrete truncated Gaussian distribution with standard deviation $\tilde{B}$, as in Definition 3.12. By Fact 3.5, the above total variation distance equals
$=\left\lVert\mathcal{D}\left(\mathbf{v}+D^{n}_{\tilde{B}}\right)-\mathcal{D}\left(\mathbf{v}+D^{n}_{\tilde{B}}+\mathbf{t}\right)\right\rVert_{TV}+2\cdot d_{\left(\mathbf{e},D^{n}_{\tilde{B}}\right)}\enspace,$ (11)
for some $d_{\left(\mathbf{e},D^{n}_{\tilde{B}}\right)}\in\mathbb{R}$ inside the range
$\bigg{[}-\left\lVert\mathcal{D}\left(D^{n}_{\tilde{B}}\right)-\mathcal{D}\left(\mathbf{e}\right)\right\rVert_{TV},\left\lVert\mathcal{D}\left(D^{n}_{\tilde{B}}\right)-\mathcal{D}\left(\mathbf{e}\right)\right\rVert_{TV}\bigg{]}\enspace,$
We next observe that the complexity of the reduction is polynomial in the input size, and then turn to proving correctness, that is, that when $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$ then the output distributions of the circuits are statistically close (completeness), and when the input is in $\prod_{\text{NO}}$, then the output distributions are statistically far (soundness). To this end, we will first bound $\left\lVert\mathcal{D}\left(D^{n}_{\tilde{B}}\right)-\mathcal{D}\left(\mathbf{e}\right)\right\rVert_{TV}$ (which in turn will bound the absolute value of $d_{\left(\mathbf{e},D^{n}_{\tilde{B}}\right)}$), and then analyze the soundness and completeness separately.